# Diffusion Models in Simulation-Based Inference: A Tutorial Review

### v1

@@ A.1.2 Case Study 2: Large-Scale ODE Benchmark
Synthetic datasets are simulated under the original experimental conditions with parameters sampled from normal priors on the parameter scale, centered at the published parameter estimates. Noise is added according to the Gaussian measurement model specified. This allows to compute the likelihood and its gradients and controlled comparison between Markov chain Monte Carlo (MCMC) and a range of diffusion based approaches. We use AMICI [frohlich2020amici] for simulation and an adaptive parallel tempering algorithm where each chain consists of a Metropolis-Hastings sampler with adaptive proposal covariance implemented in pyPESTO [schalte2023pypesto] as overall multi-chain algorithms perform better than single-chain algorithms [ballnus2017comprehensive]. Due to the high dimensions, exploring the parameter space is difficult even for the adaptive parallel tempering algorithm.
@@
To obtain a scalar comparison across methods, we rank them by the sum of ECE and NRMSE. For this setting, we combine the diffusion models with a summary network, which embeds the data before passing it through the same backbone as before. All models were tested with two distinct summary networks implemented in BayesFlow [bayesflow_2023_software]:
@@
  • And a TimeSeriesNetwork that follows the LSTNet architecture [ZhangMik2023]. It combines convolutional layers to extract local temporal features with a LSTM with 256 units that captures long-range dependencies.
@@
For sampling, we use adaptive ODE and SDE solvers. We used TSIT5 as ODE solver and the two-step adaptive solver by [Jolicoeur-MartineauLi2021] with settings as described before in Section A.1.1. As a classifier, we use a sequence of three MLPs with width of 128 and a validation split of 0.25.
@@ A.1.3 Case Study 3: Gaussian Random Fields as a Scalable SBI Benchmark
We implement the GRF prior using the FyeldGenerator routine generate_field [cadiou2022-fyeldgenerator]. This function takes as input (i) a sampling rule for a random field in Fourier space, (ii) a target power spectrum, (iii) the desired output shape, and (iv) a physical pixel size, and returns a real-valued random field on a regular grid.
@@ A.1.4 Case Study 4: Pooling Regimes in Cognitive Modeling
For this case study, we have two different priors: a flat and a hierarchical one. For each we train a diffusion model with EDM noise schedule with the standard backbone as in the first case studies. For training, we simulate 256 batches of size 128, with $R=1$ subject per training instance and $N=30$ trials per subject. Hence, as a summary network, we employ a SetTransformer [lee2019set] with a summary dimension of 16 and dropout of 0.1 to account for the exchangeability of the trials. For inference, we simulate 100 test sets of 100 subjects, each with 30 trials. To solve the reverse SDE, we employ the two-step adaptive method by [Jolicoeur-MartineauLi2021] with a maximum of 1,000 steps.
@@
For inference in the no-pooling case, we select one subject per dataset, and we assume independence across subjects. In the case of complete pooling, we assume that all subjects in one dataset are described by the same parameters [Bardenet2017talldata] and we use the compositional score over all subjects following [arruda2025compositional]. We use a mini-batch size of 3 for the compositional score, and a damping factor of $1/\text{number of subjects}$.
@@
Each subject specific parameter $\boldsymbol{\theta}^{(r)}$ is drawn from a group level distribution governed by a global mean $\boldsymbol{\mu}$ and a global standard deviation $\boldsymbol{\sigma}$. This allows subjects to differ, while introducing a hierarchical prior that ties them together [habermann2024amortized]:
@@
In this setting, we can train a global model to predict the global parameters $(\boldsymbol{\mu},\boldsymbol{\sigma},\beta)$ on individual subjects with $N$ trials. Similar to the complete pooling case, and with the same setting, we get samples of the global means and standard deviations with compositional diffusion, following [arruda2025compositional]. A local model, also a diffusion matching model trained on single individuals with $N$ trials, but with an additional condition, which are the global parameters, predicts us the per subject parameters $(\nu^{(r)},\alpha^{(r)},t_{0}^{(r)})$ conditional on the global parameters. The local model conditions on observations of a single subject and, during training, the true global parameters. For inference, we perform ancestral sampling: draw a global parameter from the group level posterior and then condition the local prediction on this global sample.

### v2

@@ A.1.2 Case Study 2: Large-Scale ODE Benchmark
Synthetic datasets are simulated under the original experimental conditions with parameters sampled from normal priors on the parameter scale, centered at the published parameter estimates. Noise is added according to the Gaussian measurement model specified. This allows to compute the likelihood and its gradients and controlled comparison between Markov chain Monte Carlo (MCMC) and a range of diffusion based approaches. We use AMICI [50] for simulation and an adaptive parallel tempering algorithm where each chain consists of a Metropolis-Hastings sampler with adaptive proposal covariance implemented in pyPESTO [148] as overall multi-chain algorithms perform better than single-chain algorithms [8]. Due to the high dimensions, exploring the parameter space is difficult even for the adaptive parallel tempering algorithm.
@@
To obtain a scalar comparison across methods, we rank them by the sum of ECE and NRMSE. For this setting, we combine the diffusion models with a summary network, which embeds the data before passing it through the same backbone as before. All models were tested with two distinct summary networks implemented in BayesFlow [140]:
@@
  • And a TimeSeriesNetwork that follows the LSTNet architecture [202]. It combines convolutional layers to extract local temporal features with a LSTM with 256 units that captures long-range dependencies.
@@
For sampling, we use adaptive ODE and SDE solvers. We used TSIT5 as ODE solver and the two-step adaptive solver by [87] with settings as described before in Section A.1.1. As a classifier, we use a sequence of two MLPs with width of 128 and 5-fold cross-validation.
@@ A.1.3 Case Study 3: Gaussian Random Fields as a Scalable SBI Benchmark
We implement the GRF prior using the FyeldGenerator routine generate_field [19]. This function takes as input (i) a sampling rule for a random field in Fourier space, (ii) a target power spectrum, (iii) the desired output shape, and (iv) a physical pixel size, and returns a real-valued random field on a regular grid.
@@ A.1.4 Case Study 4: Pooling Regimes in Cognitive Modeling
For this case study, we have two different priors: a flat and a hierarchical one. For each we train a diffusion model with EDM noise schedule with the standard backbone as in the first case studies. For training, we simulate 256 batches of size 128, with $R=1$ subject per training instance and $N=30$ trials per subject. Hence, as a summary network, we employ a SetTransformer [99] with a summary dimension of 16 and dropout of 0.1 to account for the exchangeability of the trials. For inference, we simulate 100 test sets of 100 subjects, each with 30 trials. To solve the reverse SDE, we employ the two-step adaptive method by [87] with a maximum of 1,000 steps.
@@
For inference in the no-pooling case, we select one subject per dataset, and we assume independence across subjects. In the case of complete pooling, we assume that all subjects in one dataset are described by the same parameters [12] and we use the compositional score over all subjects following [4]. We use a mini-batch size of 3 for the compositional score, and a damping factor of $1/\text{number of subjects}$.
@@
Each subject specific parameter $\boldsymbol{\theta}^{(r)}$ is drawn from a group level distribution governed by a global mean $\boldsymbol{\mu}$ and a global standard deviation $\boldsymbol{\sigma}$. This allows subjects to differ, while introducing a hierarchical prior that ties them together [65]:
@@
In this setting, we can train a global model to predict the global parameters $(\boldsymbol{\mu},\boldsymbol{\sigma},\beta)$ on individual subjects with $N$ trials. Similar to the complete pooling case, and with the same setting, we get samples of the global means and standard deviations with compositional diffusion, following [4]. A local model, also a diffusion matching model trained on single individuals with $N$ trials, but with an additional condition, which are the global parameters, predicts us the per subject parameters $(\nu^{(r)},\alpha^{(r)},t_{0}^{(r)})$ conditional on the global parameters. The local model conditions on observations of a single subject and, during training, the true global parameters. For inference, we perform ancestral sampling: draw a global parameter from the group level posterior and then condition the local prediction on this global sample.